OpenStudy (anonymous):

Solve for x. 4(e^(3x)) + 2 = 22

OpenStudy (anonymous):

First subtract 2 from both sides. Then divide both sides by 4.

OpenStudy (anonymous):

That should give you e^(3x) = 5 Take the natural log of both sides. ln e^(3x) = ln5

OpenStudy (anonymous):

4*e^3x=20 e^3x=5 3x=ln5 x=(ln5)/3

OpenStudy (anonymous):

\[x=\ln 5 \div 3\]

OpenStudy (anonymous):

This gives you 3x = ln5 Divide both sides by 3 x = (ln5)/3

OpenStudy (anonymous):

myne looks better... :) lol .. (no offence)